What are the wavelength ranges for the following?

 (a) the AM radio band (540-1600 kHz) 

maximum wavelength _______ m 

minimum wavelength _______ m

 (b) the the FM radio band (88-108 MHz) 

maximum wavelength _______ 

minimum wavelength _______  
Answer #1

here,

speed of EM waves , c = 3 * 10^8 m/s

a)

the maximum wavelength , lamda1 = c /(540 * 10^3)

lamda1 = 3 * 10^8 /540 m = 556 m

the minimum wavelength , lamda2 = c /(1600 * 10^3)

lamda2 = 3 * 10^8 /(1600 * 10^3) m = 188 m

b)

the maximum wavelength , lamda1 = c /(88 * 10^6)

lamda1 = 3 * 10^8 /(88 * 10^6) m = 3.41 m

the minimum wavelength , lamda2 = c /(108 * 10^6)

lamda2 = 3 * 10^8 /(108 * 10^6) m = 2.78 m